# Distributed Lloyd-Based Algorithm for Uncertainty-Aware Multi-Robot Under-Canopy Flocking

### v2

@@ III-C Dealing with dynamics and uncertainties
The real robot deployment requires to consider a set of additional issues. First, more complex dynamics of the robots must be applied as $\dot{p}_{i}=u_{i}$ is not directly applicable. Second, tracking error, which can be caused by external disturbances or model mismatches, generating a difference between the desired and the actual velocity, has to be taken in to account. Finally, we have to consider measurement uncertainties arising from limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [10, 2]), which steers each ii–th robot towards the centroid position $c_{\mathcal{F}_{i}}$, by constraining the trajectory to not exceed its $\mathcal{F}_{i}$ cell. By doing that, we can account for the robot dynamics without violating the safety and proximity maintenance constraints. Hence, each robot $i$ solves the following optimization problem (for the sake of clarity we discard the $i$ indexing),
$\operatorname*{\min\text{imize}}_{{x}_{k},{u}_{k}}$ $\sum^{N_{t}}_{k=0}J({x}_{k},{u}_{k},c_{\mathcal{F}}),$ (13a)
@@
$p_{k}\in\mathcal{F},\,\,\,\forall k\in(0,N_{t}),$ (13d)
${x}_{0}={x}_{\textrm{init}},$ (13e)
@@
where $k$ indicates the time index, $N_{t}$ indicates the horizon length, $J$ represents the cost function, ${x}_{k}$ and ${u}_{k}$ are the state and control inputs respectively, $f({x}_{k-1},{u}_{k-1})$ indicates the dynamics of the robot, ${x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U}$ are the state and input constraints respectively, $p_{k}\in\mathcal{F}$ constrains the future positions of the robot in its cell, and ${x}_{0}={x}_{\textrm{init}}$ is the initial condition.
@@
Both the tracking errors and measurement uncertainties can be accounted for by means of an adaptive strategy. In fact, it can be achieved by adapting the value of $\beta_{i}$ in (11) in an online fashion. An alternative approach would be reshaping the $\mathcal{F}_{i}$ cell, similarly to [10]. However, this method cannot be applied in our case, since, by following this approach, the robot may fall out of its cell due to tracking error, and once it happens, both the cell and the centroid position are no longer defined. Conversely, the value of $\beta_{i}$ defines the aggressiveness of the ii–th robot. A smaller value corresponds to a centroid position closer to the goal (and hence closer to the boundaries of the $\mathcal{F}_{i}$ cell, namely $\partial\mathcal{F}_{i}$), while larger values of $\beta_{i}$ push the centroid position as far as possible from the boundaries of the cell $\partial\mathcal{F}_{i}$. Thus, managing errors and uncertainties is achieved by selecting a value of $\beta_{i}$ such that the centroid position stays at a controlled distance from the boundary of the cell. This distance must be chosen in accordance with the maximum positional tracking error and the maximum error in the neighbors’ position estimates. More formally, it can be done by imposing a lower bound to $\beta_{i}$ in (11) equal to
$\beta^{\min}_{i}=\operatorname*{arg\,min}_{\beta_{i}}\left(\|c_{\mathcal{F}_{i}}-p_{\partial\mathcal{F}_{i}}\|-d_{u}\right)^{2},$
@@
where $p_{\partial\mathcal{F}_{i}}$ is the position of the closest point on the contour of $\mathcal{F}_{i}$, while $d_{u}$ is a threshold value that has to be tuned on the basis of the tracking errors and measurement uncertainties. For the sake of clarity, we depicted three scenarios in Figure 5. Figure 5a considers the tracking error; in this case the value of $c_{\mathcal{F}_{i}}$ is controlled to stay at least at a distance $d^{t}_{u}$ from the boundary $\partial\mathcal{F}_{i}$. We indicate with $\mathcal{Z}_{i}$ the region set where the ii–th robot is allowed to go in the next time step. Notice that the dimension of the region $\mathcal{Z}_{i}$ can be controlled by $d_{u}^{t}$. To be conservative, the value of $d_{u}^{t}$ should be determined experimentally as the maximum tracking error, and it can be set as a function of the velocity. In Figure 5b, we show the case where we have measurement uncertainties. We illustrate both the ii–th and the jj–th robots’ positions and the ellipse of uncertainty $\mathcal{E}_{i,j}$. In this case, $\mathcal{K}_{i}$ indicates the region where the centroid $c_{\mathcal{F}_{i}}$ has to be constrained to preserve safety. By relying on geometric reasoning, depending on the measurement uncertainty, when $\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|<k\sqrt{\lambda^{\max}}+\delta$, we can tune the parameter $d_{u}^{m}$ as
$d_{u}^{m}\geq\left\|\left(1-\frac{1}{\epsilon_{p}}\right)\|p_{i}-p_{j}\|-\left({k\sqrt{\lambda^{\max}}}+\delta\right)\right\|+\delta,$ (14)
@@
where $\lambda^{\max}$ is the maximum eigenvalue of the covariance matrix associated with the measurement and $k$ is a scale factor that defines the confidence ellipse (e.g., $k\approx 2.795$ for a confidence level of $95\%$).
@@
From a practical point of view, the value of $d_{u}^{m}$ may be set empirically to a constant that satisfies inequality (14) in the worst-case expected scenario. In real-world equipments the uncertainties associated with the measurements are far from being constant values. Hence, the user can set $d_{u}^{m}$ dynamically, based on the current measurement uncertainties, which are reported by the equipment together with the measurement means. One method to implement this, is to use the restraining method from [31]. This technique offsets the set-points of the robots from an ideal target location according to the probability distributions of relative pose observations.
On the other hand, to preserve proximity maintenance, if $\Gamma_{ij}-\|p_{i}-p_{j}\|\approx k\sqrt{\lambda^{\max}}$, we need to set $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$. Since this value may be quite high in practice, it is sufficient to impose that if the mean value of the neighbor position is out of the proximity area, then we can re-project its position to the closest point inside the proximity area. In this way, the proximity to the neighbors is effectively maintained.
Finally, we show the combination of tracking error and measurement uncertainties, where $d_{u}=d_{u}^{m}+d_{u}^{t}$ in Figure 5c.
@@
Remark 2 (Convergence for the adaptive solution).
By relying on the adaptive solution, while safety and proximity maintenance have been taken into account, we lost the convergence guarantees. In fact, by relying on the adaptive strategy that accounts for tracking error and measurement uncertainties, a new sufficient condition is introduced, which strictly depends on $d_{u}$. In particular, given the minimum distance between the obstacles in the environment being $d^{\min}_{o}$, in order to ensure the passage of the robot between two obstacles, it can be shown through geometrical reasoning, that we need to have $0\leq d_{u}<\frac{d^{\min}_{o}}{2}-(\delta_{i}+\delta^{o})$. A situation where the cell is too small to allow the centroid to be at the required distance from the boundary of the cell can also occur. However, we claim that this is not an issue for safety nor for proximity maintenance, since in these cases the robot velocity approaches zero. Such situation is transitory and may occur in very crowded conditions or when the robot has to maintain proximity with two or more robots that are close to the maximum allowed distance $\Gamma_{ij}$.
@@ IV Simulation results
In this section we show through simulations, the scalability, repeatability and robustness of our approach, and we provide a comparison with a state-of-the-art algorithm. In particular, we report a comparison with PACNav [7], an algorithm that does not rely on communication network, and shares similar features with the proposed algorithm (refer to Table I). Notice that comparing with algorithm such as [5, 4, 6] is out of scope of this paper. Our method does not perform better than algorithms that can share information, such as future trajectories, among the robots. Despite [7] performs well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios. In addition, crucial requirements such as safety and proximity maintenance depend strongly on fine parameter tuning. This is a common issue in the state-of-the-art algorithms, which is overcome in our solution, where the parameter selection affects only the performance of the mission, while, safety, proximity maintenance and convergence are met if they are selected according to the theoretical findings. We perform the comparison in the MRS simulator [32] in the same environment and same model dynamics for both algorithms. We simulate three scenarios with different number of robots and with an increasingly complex obstacle configurations. For each scenario, we perform $10$ runs for statistical evidence. The metrics that we took into account are the time required for all the robots to travel $42$ meters in the goal direction and the success rate in terms of mutual collision avoidance and obstacle avoidance. In Table II, we report the results obtained from PACNav [7] and from our approach. We considered the following scenario: $N=[4,9,9]$, $\delta_{i}=[0.25,0.5,0.5]$ (m), $\delta_{o}=[0.3,0.75,0.75]$ (m), and we selected the parameters as follows for all the three scenarios $r_{s,i}=4.5$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=0.5$, $\Gamma_{ij}=8$ (m), $d_{u}=0.8$ (m). The adjacency matrix which defines the proximity maintenance constraints were selected as,
$A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right],\text{ and}\quad A_{d}=\left[\begin{smallmatrix}0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 0&1&0&0&0&1&0&0&0\\ 1&0&0&0&1&0&1&0&0\\ 0&1&0&1&0&1&0&1&0\\ 0&0&1&0&1&0&0&0&1\\ 0&0&0&1&0&0&0&1&0\\ 0&0&0&0&1&0&1&0&1\\ 0&0&0&0&0&1&0&1&0\\ \end{smallmatrix}\right].$
@@
For our approach we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$. The parameter selection for PACNav [7] has been done by trial and error, on the basis of the environment and the number of robots. Besides that, as seen in Table II, the approach does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
@@
Table II: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in 33 different scenarios, comparing PACNav [7] with our method in two different parameter configurations.
For further simulation results, please refer to the attached multimedia material, where we include simulations with up to $16$ UAVs in crowded conditions. Notice that even if the sufficient condition in Proposition 1 is not respected, the robots can still safely reach the goal in environments with $56\%$ of free space, as shown in multimedia material.
@@ V Experimental Results
The experimental deployment is conducted in the forest, where we do not have access to the GNSS signal. The goal positions are set to be $40$ (m) ahead with respect to the initial position of each robot, notice that we do not need to have a common reference frame. However, we need to satisfy the requirement stated in Proposition 1, i.e., $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right]$.
@@
In Figure 8-(a) we show the trajectory of UAV $2$, the positions of the other robots measured by UVDAR system, and the map built by UAV $2$ by using Hector SLAM during the mission. In Figure 6, we depict the distances between UAV $2$ and the other robots according to the UVDAR measurements. In Table III, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission. To verify the success of the mission, we report in Figure 8-(b) the actual trajectories for all the 4 robots in the mission, while in Figure 7 we show the distances between the UAVs during the mission. As we expected the safety constraints and the flocking constraints (according to the adjancency matrix $A_{d}$) are satisfied.
@@
To show reliability and robustness to parameter changes, we report in Table IV the results obtained from other four experiments in the forest, with $N=3$ UAVs. Since the system is intrinsically chaotic it does not make sense to compare the results obtained, even if, as we can expect, a relation between $\beta^{D}$ and the time to reach the goal can be inspected from Table IV. Nevertheless, through the table we want to show the repeatability of the experiments and the robustness to parameter changes. The parameters for each experiment were set as follows: #\#experiment =[1,2,3,4]=[1,2,3,4], $N=[3,3,3,3]$, $\delta_{i}=[0.3,0.3,0.3,0.3]$ (m), $\delta^{o}=[0.05,0.05,0.05,0.05]$ (m) $r_{s,i}=[6,6,6,6]$ (m), $d_{1}=d_{3}=[0.5,0.5,0.5,0.5]$ (m), $d_{2}=d_{4}=[1,1,1,1]$ (m), $\beta_{i}^{D}=[2.0,2.0,1.0,0.6]$, $\Gamma_{ij}=[12,10,10,10]$ (m), $d_{u}=[0.75,0.75,0.75,0.75]$ (m), $\epsilon_{p}=\epsilon_{o}=[1.052,1.052,1.010,1.052]$, and the adjacency matrix that defines the flocking constraints is the same for all the four experiments, $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right].$
@@
Table III: Average speed, maximum speed, and time to reach the goal of each UAV in the experiment from Figure 8.
Figure 7: Distances between UAVs computed based on odometries transformed to a common reference frame obtained through alignment of maps of individual UAVs in a post-prossessing phase.
@@
Table IV: Comparison of average speed, maximum speed, and time to reach the goal in four different experiments in the forest, showing repeatability and robustness to parameters changes.

### v3

@@ III-C Dealing with dynamics and uncertainties
The robot deployment requires to consider a set of additional issues. First, more complex dynamics of the robots must be applied as $\dot{p}_{i}=u_{i}$ is not directly applicable. As a consequence, tracking error, which is caused by model mismatch and external disturbances, generating a difference between the desired and the actual velocity, has to be taken into account. Finally, we have to consider measurement uncertainties arising from limitations of the perception system providing an estimate of the positions of other robots and obstacles with a certain error.
To account for the dynamic constraints of the robots, we synthesize the control inputs through an MPC formulation (similarly to [32, 9]). Hence, each robot $i$ solves the following optimization problem (for the sake of clarity we discard the $i$ indexing),
$\operatorname*{\text{minimize}}_{{x}_{k},{u}_{k}}$ $\sum^{N_{t}}_{k=0}J({x}_{k},{u}_{k},c_{\mathcal{F}}),$ (13a)
@@
${x}_{0}={x}_{\textrm{init}},$ (13d)
@@
where $k$ indicates the time index, $N_{t}$ indicates the horizon length, $J$ represents the cost function, ${x}_{k}=[p_{k}^{\top},\dot{p}_{k}^{\top},\ddot{p}_{k}^{\top}]^{\top}$ and ${u}_{k}$ are the state and control inputs respectively, $f({x}_{k-1},{u}_{k-1})$ indicates the dynamics of the robot, ${x}_{k}\in\mathcal{X},\,{u}_{k}\in\mathcal{U}$ are the state and input constraints respectively, and ${x}_{0}={x}_{\textrm{init}}$ is the initial condition.
@@
Both the tracking errors and measurement uncertainties can be accounted for by means of an adaptive strategy. Specifically, it is achieved by adapting the value of $\beta_{i}$ in (11) in an online fashion. An alternative approach would be reshaping the $\mathcal{F}_{i}$ cell, similarly to [32]. However, this method cannot be applied in our case, since, by following this approach, the robot may fall out of its cell due to tracking error, and once it happens, both the cell and the centroid position are no longer defined. Conversely, the value of $\beta_{i}$ defines the aggressiveness of the ii–th robot. A smaller value corresponds to a centroid position closer to the goal (and hence closer to the boundaries of the $\mathcal{F}_{i}$ cell, namely $\partial\mathcal{F}_{i}$), while larger values of $\beta_{i}$ push the centroid position as far as possible from the boundaries of the cell $\partial\mathcal{F}_{i}$. Thus, managing errors and uncertainties is achieved by selecting a value of $\beta_{i}$ such that the centroid position stays at a controlled distance from the boundary of the cell. This distance must be chosen in accordance with the maximum positional tracking error and the maximum error in the neighbors’ position estimates. More formally, it can be done by imposing a lower bound to $\beta_{i}$ in (11) equal to
$\beta^{\min}_{i}=\operatorname*{arg\,min}_{\beta_{i}}\left(\|c_{\mathcal{F}_{i}}-p_{\partial\mathcal{F}_{i}}\|-d_{u}\right)^{2},$ (14)
@@
where $p_{\partial\mathcal{F}_{i}}$ is the position of the closest point to $c_{\mathcal{F}_{i}}$ on the boundary of $\mathcal{F}_{i}$, while $d_{u}$ is a threshold value that has to be tuned on the basis of the tracking errors and measurement uncertainties. For the sake of clarity, we depicted three scenarios in Figure 3. Figure 3a considers the tracking error; in this case the value of $c_{\mathcal{F}_{i}}$ is controlled to stay at least at a distance $d^{t}_{u}$ from the boundary $\partial\mathcal{F}_{i}$. We indicate with $\mathcal{Z}_{i}$ the region set where the ii–th robot is allowed to go in the next time step. Notice that the dimension of the region $\mathcal{Z}_{i}$ can be controlled by $d_{u}^{t}$. To be conservative, the value of $d_{u}^{t}$ should be determined empirically as the maximum tracking error, and it can be set as a function of the velocity. In Figure 3b, we show the case where we have measurement uncertainties. We illustrate both the ii–th and the jj–th robots’ positions and the ellipse of uncertainty $\mathcal{E}_{i,j}$. In this case, $\mathcal{K}_{i}$ indicates the region where the centroid $c_{\mathcal{F}_{i}}$ has to be constrained to preserve safety. Let us consider the case where $\epsilon_{p(o)}=1$. By relying on geometric reasoning (refer to Figure 3b), depending on the inter-robot distance measurement uncertainty, we can select the parameter $d_{u}^{m}$ as
$d_{u}^{m}\geq k\sqrt{\lambda^{\max}}+\delta_{j}+\delta_{i},$ (15)
@@
where $\lambda^{\max}$ is the maximum eigenvalue of the covariance matrix associated with the distance measurement and $k$ is a scale factor that defines the confidence ellipse.
@@
From a practical point of view, the value of $d_{u}^{m}$ may be set empirically to a constant that satisfies inequality (15) in the worst expected scenario. In real-world equipments the uncertainties associated with the measurements are far from being constant values. Hence, the user can set $d_{u}^{m}$ dynamically, based on the current measurement uncertainties, which are reported by the equipment together with the measurement means. One method to implement this, is to use the restraining method from [27]. This technique offsets the set-points of the robots from an ideal target location according to the probability distributions of relative pose observations.
On the other hand, to preserve proximity maintenance, a sufficient condition is that $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$. Finally, the tracking error and measurement uncertainties are combined as $d_{u}=d_{u}^{m}+d_{u}^{t}$ in Figure 3c.
Hence, by imposing the control $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), with $\mathcal{F}_{i}=\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}\cap\mathcal{M}_{i}$ and assuming a maximum uncertainty in the estimation of the other robots’ distances equal to $k\sqrt{\lambda^{\max}}$, by selecting $d_{u}^{m}\geq k\sqrt{\lambda^{\max}}$, and assuming a bounded maximum tracking error less or equal than $d^{t}_{u}$, both safety and proximity maintenance are preserved by selecting $\beta_{i}^{\min}$ according to (14) for all $i=1,\ldots,N$. Specifically, let us choose $d_{u}^{m}$ such that it satisfies (15) and $d_{u}=d^{t}_{u}+d_{u}^{m}$. By enforcing a lower bound $\beta_{i}^{\min}$ (14), we set a minimum distance $d_{u}$ between the centroid position $c_{\mathcal{F}_{i}}$ and the boundary of the cell $\partial\mathcal{F}_{i}$. Due to the definition of $d_{u}$, this margin explicitly accounts for both estimation uncertainty and tracking error, thereby ensuring safety and proximity maintenance for all admissible realizations of the neighbors’ positions within the confidence ellipse. More intuitively, the centroid is constrained to remain within a region that preserves safety and proximity maintenance by accounting for tracking error and all possible neighbors’ positions consistent with the uncertainty bounds. If there exists no point $q\in\mathcal{F}_{i}$ such that $\text{dist}(q,\partial\mathcal{F}_{i})\geq d_{u}$, according to (14), the centroid position $c_{\mathcal{F}_{i}}$, is the point that maximizes the distance from $\partial\mathcal{F}_{i}$. Under this condition, strict guarantees of safety and proximity maintenance cannot be ensured for all possible uncertainty realizations. Nevertheless, in this case the robot adopts the most conservative behavior, placing the centroid at the location that maximizes boundary clearance. The guarantees provided by the proposed conditions are therefore sufficient and hold whenever a feasible margin exists; they rely on worst-case realizations within the assumed uncertainty bounds and bounded tracking errors.
@@ IV Simulation results
In this section, we first show through simulations the scalability, repeatability and robustness of our approach in two different scenarios, with and without accounting for uncertainties and tracking errors. Then we report a comparison with PACNav [2], the algorithm does not rely on communication network, and shares similar features with the proposed algorithm (refer to Table I). Notice that comparing with algorithm such as [30, 29, 31] is out of scope of this paper. In most of the scenarios, our method does not perform better than algorithms that can share information, such as future trajectories, among the robots. Despite [2] performs well in relatively simple environments and with a limited number of robots, it lacks reliability in more complex scenarios. In addition, crucial requirements such as safety and proximity maintenance depend strongly on fine parameter tuning. This is a common issue in the state-of-the-art algorithms, which is overcome in our solution, where the parameter selection affects only the performance of the mission, while, safety and proximity maintenance are met if they are selected according to the theoretical findings. All the simulations are performed by relying on the MRS simulator [3], which account for full rigid-body UAV dynamics.
In the first simulations, we consider different set of parameters $d_{u}=[0,1.0]$ (m) in two different scenarios, with $N=[9,16]$, $\delta_{i}=0.2$ (m), $\epsilon=1.0$, and randomly generated circular obstacles with $\delta_{o}=0.15$ (m) (see Figure 4). We select the parameters as follows for all the simulations $r_{s,i}=5.0$ (m), $d_{1}=d_{2}=d_{3}=d_{4}=1$ (m), $\beta_{i}^{D}=0.15$, $\Gamma_{ij}=10$ (m), while the adjacency matrices defining proximity maintenance constraints are defined as grid matrices, as depicted in Figure 4 (red lines). We simulated UVDAR noise as a bounded radial random measurement error with magnitude not exceeding $0.8$ (m). In Table II, we report the quantitative results in terms of average time to reach the goal and success rate over a total of $40$ runs. As expected, setting $d_{u}=0$ (m) leads to better performance in terms of time to reach the goal; however, it does not ensure collision avoidance between the robots. In contrast, selecting $d_{u}$ according to the theoretical findings results in safe behavior and preserves the proximity maintenance constraints. Figure 4 illustrates the two scenarios with $N=[9,16]$ robots, and shows the trajectories generated for one instance with $d_{u}=1.0$ (m). Figure 5 illustrates the distance between each robot in the two scenarios. The colored lines indicate whether the proximity maintenance constraint is active. To further validate the robustness of the proposed approach, we additionally performed 200 simulations across four different parameter sets in the same scenario depicted in Figure 4a, consistently achieving the goal preserving safety and the proximity maintenance constraints.
Refer to caption Refer to caption
(a) (b)
Figure 4: Simulation results with $N=9$ and $N=16$ robots, with $d_{u}=1.0$ (m). The images represent example scenarios used for generation of results presented in Table II. Colored circles represent the robots’ positions at different time $t$. Red lines represent the proximity maintenance constraints between the robots. Black dots are the obstacles in the scene.
Figure 5: Distances between UAVs in the simulations in Figure 4a and 4b. The flocking and safety constraints are in magenta. Colored lines indicate the presence of an active flocking constraint between the two robots.
$d_{u}=0\penalty 10000\ (m)$ $d_{u}=1.0\penalty 10000\ (m)$
N avg. time (s) SR avg. time (s) SR
9 32.71 0.50 54.91 1.00
16 34.94 0.40 64.53 1.00
Table II: The simulation results obtained in terms of time to reach the goal and success rate (SR) for $10$ instances in 22 different scenarios with $N=[9,16]$. We report the results with different parameter selection i.e., $d_{u}=0$ (m) and $d_{u}=1.0$ (m).
For the comparison with PACNav [2], we simulate three scenarios with different number of robots and a varying number of randomly generated circular obstacles. In the simulation, we do not inject noise in the measurements, nevertheless tracking error is present. For each scenario, we perform $10$ runs for statistical evidence. The metrics that we took into account are the time required for all the robots to travel $42$ meters in the goal direction and the success rate in terms of mutual collision avoidance and obstacle avoidance. In Table III, we report the results obtained from PACNav [2] and from our approach. We considered the following scenarios: $N=[4,9,9]$, $\delta_{i}=[0.25,0.5,0.5]$ (m), $\delta_{o}=[0.3,0.75,0.75]$ (m), and we select the parameters as follows for all the three scenarios $r_{s,i}=4.5$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=0.5$, $\Gamma_{ij}=8$ (m), $d_{u}=0.8$ (m). The adjacency matrices defining the proximity maintenance constraints were selected as grid adjacency matrices.
@@
For our approach, we considered two different values for the parameters $\epsilon=\epsilon_{p}=\epsilon_{o}$. The parameter selection for PACNav [2] has been done by trial and error approach, on the basis of the environment and the number of robots. Besides that, as seen in Table III, PACNav does not scale in terms of performance (convergence, safety, and proximity maintenance) when the number of robots grows, due to the complex superimposition of forces given by the surroundings.
@@
Table III: The simulation results obtained in terms of time to reach the goal and success rate for $10$ instances in 33 different scenarios, comparing PACNav [2] with our method in two different parameter configurations.
For further simulation results, please refer to the multimedia material.
@@ V Experimental Results
The experimental deployment is conducted in the forest, where we do not have access to the GNSS signal. The goal positions are set to be $40$ (m) ahead with respect to the initial position of each robot, notice that we do not need to have a common reference frame. We choose to adopt a fixed value of $d_{u}$ in the experiments. Due to the high measurements’ uncertainties the value of $d_{u}$ did not satisfy (15) all the time. In this regard, two practical solutions were adopted in the experiments: (i) We discard measurements with values of $\lambda^{\max}>15$, and (ii) if the measured distance between $i$-th and $j$-th robot with $j\in\bar{\mathcal{N}_{i}}$ is greater than $\Gamma_{ij}$, we re-project the position of the $j$-th robot to be at a distance less than $\Gamma_{ij}$ from the $i$-th robot, in order to ensure the existence of a nonempty set $\mathcal{F}_{i}$. These two strategies help to practically manage large estimation errors that exceed the theoretical threshold e.g., temporary failures in the estimation of distance to other robot’s due to tree occlusions.
We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right]$.
@@
In Figure 6-(a) we show the trajectory of UAV $2$, the positions of the other robots measured by UVDAR system, and the map built by UAV $2$ by using Hector SLAM during the mission. In Figure 7, we depict the distances between UAV $2$ and the other robots according to the UVDAR measurements. In Table IV, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission. To verify the success of the mission, we report in Figure 6-(b) the actual trajectories for all the 4 robots in the mission, while in Figure 8 we show the distances between the UAVs during the mission. The safety constraints and the flocking constraints (according to the adjancency matrix $A_{d}$) were satisfied during the entire mission.
@@
To show reliability and robustness to parameter changes, we report in Table V the results obtained from other four experiments in the forest, with $N=3$ UAVs. Since the system is intrinsically chaotic it does not make sense to compare the results obtained, even if, as we can expect, a relation between $\beta^{D}$ and the time to reach the goal can be inspected from Table V. Nevertheless, through the table we want to show the repeatability of the experiments and the robustness to parameter changes. The parameters for each experiment were set as follows: #\#experiment =[1,2,3,4]=[1,2,3,4], $N=[3,3,3,3]$, $\delta_{i}=[0.3,0.3,0.3,0.3]$ (m), $\delta^{o}=[0.05,0.05,0.05,0.05]$ (m) $r_{s,i}=[6,6,6,6]$ (m), $d_{1}=d_{3}=[0.5,0.5,0.5,0.5]$ (m), $d_{2}=d_{4}=[1,1,1,1]$ (m), $\beta_{i}^{D}=[2.0,2.0,1.0,0.6]$, $\Gamma_{ij}=[12,10,10,10]$ (m), $d_{u}=[0.75,0.75,0.75,0.75]$ (m), $\epsilon_{p}=\epsilon_{o}=[1.052,1.052,1.010,1.052]$, and the adjacency matrix that defines the flocking constraints is the same for all the four experiments, $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right].$
@@
Table IV: Average speed, maximum speed, and time to reach the goal of each UAV in the experiment from Figure 6.
@@
Figure 8: Distances between UAVs computed based on odometries transformed to a common reference frame obtained through alignment of maps of individual UAVs in a post-prossessing phase. The flocking and safety constraints are in magenta. According to $A_{d}$ the flocking constraints are not active for the pair of robots 22-44 and 11-33.
@@
Table V: Comparison of average speed, maximum speed, and time to reach the goal in four different experiments in the forest, showing repeatability and robustness to parameters changes.